# Ctrl-U: Robust Conditional Image Generation via Uncertainty-aware Reward Modeling

### v1

@@ Abstract
In this paper, we focus on the task of conditional image generation, where an image is synthesized according to user instructions. The critical challenge underpinning this task is ensuring both the fidelity of the generated images and their semantic alignment with the provided conditions. To tackle this issue, previous studies have employed supervised perceptual losses derived from pre-trained models, i.e., reward models, to enforce alignment between the condition and the generated result. However, we observe one inherent shortcoming: considering the diversity of synthesized images, the reward model usually provides inaccurate feedback when encountering newly generated data, which can undermine the training process. To address this limitation, we propose an uncertainty-aware reward modeling, called Ctrl-U, including uncertainty estimation and uncertainty-aware regularization, designed to reduce the adverse effects of imprecise feedback from the reward model. Given the inherent cognitive uncertainty within reward models, even images generated under identical conditions often result in a relatively large discrepancy in reward loss. Inspired by the observation, we explicitly leverage such prediction variance as an uncertainty indicator. Based on the uncertainty estimation, we regularize the model training by adaptively rectifying the reward. In particular, rewards with lower uncertainty receive higher loss weights, while those with higher uncertainty are given reduced weights to allow for larger variability. The proposed uncertainty regularization facilitates reward fine-tuning through consistency construction. Extensive experiments validate the effectiveness of our methodology in improving the controllability and generation quality, as well as its scalability across diverse conditional scenarios, including segmentation mask, edge, and depth conditions. Code, models, and demo will soon be available at this https URL.
@@ 1 Introduction
Driven by the emergence of large-scale image-text datasets (Schuhmann et al., 2021; Changpinyo et al., 2021; Schuhmann et al., 2022) and the development of diffusion models (Dhariwal & Nichol, 2021; Nichol et al., 2021; Rombach et al., 2022), text-to-image (T2I) diffusion models (Ramesh et al., 2021; Saharia et al., 2022) have become fundamental in the field of controllable visual generation. These models excel at producing realistic, high-quality images that accurately reflect the descriptions provided in natural language. However, due to their inherent properties, text-based conditions often struggle to convey the detailed controls required for final generation results efficiently. This limitation becomes especially evident in certain scenarios where text prompts alone can not capture all necessary details, such as when depicting unique artistic styles or accurately rendering complex scenes. To this end, alongside text descriptions, a substantial body of research focuses on integrating novel conditional controls, such as user-drawn sketches and semantic masks, into T2I diffusion models (Qin et al., 2024; Ye et al., 2023; Zhang et al., 2023a; Mou et al., 2024). Despite investigations into the controllability of T2I diffusion models and their expanded applications (Gao et al., 2024; Li et al., 2024a; Chen et al., 2024a), achieving precise and fine-grained control remains challenging. The primary issue lies in ensuring both the fidelity of the generated images and their semantic alignment with the provided conditions. To tackle this issue, some efforts, e.g., ControlNet++ (Li et al., 2024b) attempt to employ a pre-trained reward model to extract the corresponding condition from the generated images and enforce alignment between the specified condition and the output. However, we observe that the reward model inevitably produces inaccurate feedback (see Fig. 1). As the timestep $t$ increases, and the generation diversity also increases. Some newly generated data reflects the correct condition but has not been "seen" by the reward model during training, resulting in inaccurate feedback. For instance, although the denoised image at $t=0$ aligns well with the input condition, there are still some mIoU errors according to the reward model. If we do not rectify such misleading rewards, enforcing alignment between the provided conditions and the inaccurate predictions on newly generated data will consistently compromise the training process.
@@ 2 Related Works
Conditional Generation. Recently, diffusion probabilistic models (Ho et al., 2020; Song et al., 2021a) have become an important cornerstone in general image generation. As the quality of generated images improves, a critical challenge remains, i.e., achieving precise control over these generative models to adequately meet the intricate and varied demands of real-world applications. Motivated by the development of guidance mechanism (Song et al., 2021b; Ho & Salimans, 2022), T2I diffusion models such as GLIDE (Nichol et al., 2021), Imagen (Saharia et al., 2022), DALL·E 2 (Ramesh et al., 2022) excel at modeling fine-grained structures and texture details. To facilitate the training of diffusion models on limited resources, Latent Diffusion Model (LDM) (Rombach et al., 2022) maps the diffusion process from pixel space to the latent space. Building upon the foundation of LDM framework, Stable Diffusion (SD) exhibits exceptional capabilities in T2I generation. Despite the astonishing capabilities of T2I diffusion models, language, with its sparse and high-level semantic nature, is unsuitable for processing intricate and low-level control images, e.g., depth maps. To achieve conditional control in T2I diffusion models, researchers are exploring the integration of various control signals with text descriptions. ControlNet (Zhang et al., 2023a) integrates image-based conditions by incorporating an additional encoder copy into frozen T2I diffusion models via zero convolutions. Similarly, GLIGEN (Li et al., 2023) and T2I-Adapter (Mou et al., 2024) propose utilizing independent adapters (or extra modules) to synchronize internal knowledge within T2I models with external conditions. Furthermore, Cocktail (Hu et al., 2023) integrates multi-modal control information within a single diffusion model to yield more high-quality outputs. With the rapid development of large language models, some recent works have explored prompt engineering for regulated generation. For instance, ReCo (Zhang et al., 2023b) specifies text descriptions of different regions to model the fine-grained structure and aesthetic features. Taking one step further, Control-GPT (Zhang et al., 2023b) applies GPT-4 (OpenAI, 2023) as the sketch generator to incorporate control signals. However, one drawback is that their generated images usually deviate from input conditions. In this work, we take a closer look at denoised results and explicitly introduce the uncertainty to adaptively impose consistency constraints.
@@
Uncertainty-aware learning. As data-driven technologies rapidly advance, the imperative for model reliability grows. Measuring the "confidence" of predictions has been a long-standing problem. As a result, researchers are progressively focusing on uncertainty as a viable solution.  Kendall & Gal (2017) categorize uncertainty into two primary types: epistemic uncertainty and aleatoric uncertainty. The former epistemic uncertainty denotes model uncertainty, reflecting variations in model weights even when trained on the same dataset. The representative approach in this direction is Bayesian networks (Jensen & Nielsen, 2007; Neal, 2012), which focuses on learning the distribution of weights and estimates uncertainty by assessing the distribution variance. Similarly, Monte Carlo Dropout (Gal & Ghahramani, 2016) is proposed to effectively simulate variational Bayesian approximation by randomly using varying dropout masks. Another line of research on epistemic uncertainty (Raghu et al., 2019; Nandy et al., 2020; Lee & AlRegib, 2020) adapts the original model to measure prediction uncertainty directly. However, incorporating this additional uncertainty estimation typically compromises both the training efficiency and predictive accuracy of the model. In addition, ensemble methods (Lakshminarayanan et al., 2017; Malinin et al., 2019; Wenzel et al., 2020) combine various deterministic models in the prediction process to improve prediction accuracy, but is constrained by the computational burden associated with operating multiple independent networks and the requisite diversity across ensemble models. Another family of research tackles the inherent uncertainty in data, arising from noise interference or ambiguous annotations. This methodology for addressing uncertainty has been applied across various fields: such as image segmentation (Zheng & Yang, 2021), image retrieval (Chen et al., 2024b), image classification (Litrico et al., 2023), person re-identification (Zhang et al., 2022), and 3D object detection (Zhang et al., 2024). In particular,  Dou et al. (2022) effectively simulates uncertainty by utilizing model prediction variance to inject noise into the latent space. In a manner akin to (Kendall & Gal, 2017), a dynamic uncertainty-based loss is introduced to enhance training stability. Unlike existing uncertainty-based works, we explicitly introduce noise at varying intensities into the diffusion model to enable the quantification of uncertainty. It is worth noting that we explore the utilization of uncertainty in the untapped conditional image generation task, preserving the original diffusion learning and mitigating the adverse effects of inaccurate feedback from the reward model.
@@ 3.1 Uncertainty Estimation
$z_{1}=\sqrt{\bar{\alpha}_{t_{1}}}z_{0}+\sqrt{1-\bar{\alpha}_{t_{1}}}\epsilon,\quad z_{2}=\sqrt{\bar{\alpha}_{t_{2}}}z_{0}+\sqrt{1-\bar{\alpha}_{t_{2}}}\epsilon,\quad\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
Following the ControlNet pipeline (Zhang et al., 2023b), we further fuse text condition $f_{t}$ and image condition $f_{c}$ to predict the added noise. After removing the predicted noise, we obtain the recovered latent, i.e., $\hat{z}_{0}^{1}$ and $\hat{z}_{0}^{2}$. Then, given the latent $\hat{z}_{0}^{1}$ and $\hat{z}_{0}^{2}$, the pretrained decoder is to reconstruct the input image as $\hat{x}_{0}^{1}$ and $\hat{x}_{0}^{2}$, respectively.
Figure 2: A brief overview of our pipeline. Here, we take the segmentation mask as a conditional generation example. (a) Conditional Generation. Given text, source image $x_{0}$, and the conditional control $c$, we extract feature $z_{0}$, $f_{t}$, $f_{c}$, respectively. Then, we fine-tune the Diffusion model to generate two intermediate features for the image decoder. (b) Uncertainty Learning. Given the two features, we decode the two images, i.e., $\hat{x}_{0}^{1}$ and $\hat{x}_{0}^{2}$. Then we apply the reward model to obtain the two layout predictions $\hat{c}_{1}$ and $\hat{c}_{2}$. We leverage the prediction discrepancy as the uncertainty indicator to rectify the original reward loss.
To align the condition within the generated images, we follow the existing works and apply an off-the-shelf reward model $D$ to quantify the consistency between the input condition $c$ and the corresponding output condition of the generated images. However, as shown in Fig. 1(b), the reward model usually contains inaccurate feedback, even if the visual layout has already aligned with the condition. If we consistently apply strong supervision on the consistency between inaccurate rewards and conditions predicted on the newly generated data, the backpropagation of incorrect gradients will significantly compromise the model. To estimate inaccrate rewards, we explicitly leverage two diffusion forwards for the identical input conditions. We compare the reward discrepancy between extracted conditions $\hat{c}_{1},\hat{c}_{2}$ from generated images, which can be considered as a reward indicator at the current timestep. For the segmentation map condition, considering the probability output, we quantify uncertainty by calculating the KL-divergence between the extracted conditions of two generations at the pixel level as:
@@ 3.2 Uncertainty Regularization
The existing controllability modeling methods (Li et al., 2024b) usually adopt consistency loss between the input condition and the extracted condition, which can be regarded as a pixel-wise supervision. For example, when using segmentation mask as input condition, $\mathcal{L}^{c}$ can be defined as a per-pixel cross-entropy loss as:
@@
$\mu=\begin{cases}\mu_{0}&\text{if }t\leq t_{\text{thre}}\\ 0&\text{if }t\textgreater t_{\text{thre}}\end{cases}$ (7)
where $\mu_{0}$ is the consistency weight. If $t\textgreater t_{\text{thre}}$, we only consider the diffusion optimization as existing works (Zhao et al., 2024a). if $t\leq t_{\text{thre}}$, we incorporate uncertainty-aware reward modeling.
Discussion. 1). Why set timestep threshold $t_{\text{thre}}$ for $\mu$? Considering that when $t$ is large, $z_{t}$ approaches the random noise $\epsilon$, leading to more diverse recovered $\hat{x}_{0}$ (see Fig. 1). In such cases, it becomes difficult to ensure that two outputs with identical conditions remain close, which is necessary for our uncertainty estimation, as it requires visually similar inputs to evaluate the reward model. In practise, we, therefore, empirically set a threshold to avoid the large timesteps for the reward modeling. We have conducted the experiment on the choice of $t_{\text{thre}}$ (see Section 4.3). 2). What is the advantage of uncertainty regularization in the reward feedback? The advantage of uncertainty regularization in the reward feedback is that it helps to mitigate the adverse effects of imprecise feedback from the reward model. Specifically, by incorporating uncertainty estimation, the proposed method, Ctrl-U, can adaptively adjust the weights of the rewards during training. Rewards with lower uncertainty are given higher loss weights, ensuring that more reliable feedback has a stronger influence on the training process. Conversely, rewards with higher uncertainty are given reduced weights, allowing for greater variability and preventing the model from being overly influenced by potentially inaccurate feedback. This adaptive rectification of rewards based on their uncertainty improves the consistency and reliability of the training process. As a result, the model becomes more robust to the diverse and sometimes unpredictable nature of newly generated data, leading to better controllability and generation quality.

### v3

@@ Abstract
In this paper, we focus on the task of conditional image generation, where an image is synthesized according to user instructions. The critical challenge underpinning this task is ensuring both the fidelity of the generated images and their semantic alignment with the provided conditions. To tackle this issue, previous studies have employed supervised perceptual losses derived from pre-trained models, i.e., reward models, to enforce alignment between the condition and the generated result. However, we observe one inherent shortcoming: considering the diversity of synthesized images, the reward model usually provides inaccurate feedback when encountering newly generated data, which can undermine the training process. To address this limitation, we propose an uncertainty-aware reward modeling, called Ctrl-U, including uncertainty estimation and uncertainty-aware regularization, designed to reduce the adverse effects of imprecise feedback from the reward model. Given the inherent cognitive uncertainty within reward models, even images generated under identical conditions often result in a relatively large discrepancy in reward loss. Inspired by the observation, we explicitly leverage such prediction variance as an uncertainty indicator. Based on the uncertainty estimation, we regularize the model training by adaptively rectifying the reward. In particular, rewards with lower uncertainty receive higher loss weights, while those with higher uncertainty are given reduced weights to allow for larger variability. The proposed uncertainty regularization facilitates reward fine-tuning through consistency construction. Extensive experiments validate the effectiveness of our methodology in improving the controllability and generation quality, as well as its scalability across diverse conditional scenarios, including segmentation mask, edge, and depth conditions. Codes are publicly available at https://grenoble-zhang.github.io/Ctrl-U.
@@ 1 Introduction
Driven by the emergence of large-scale image-text datasets (Schuhmann et al., 2021; Changpinyo et al., 2021; Schuhmann et al., 2022) and the development of diffusion models (Dhariwal & Nichol, 2021; Nichol et al., 2021; Rombach et al., 2022), text-to-image (T2I) diffusion models (Ramesh et al., 2021; Saharia et al., 2022) have become fundamental in the field of controllable visual generation. These models excel at producing realistic, high-quality images that accurately reflect the descriptions provided in natural language. However, due to their inherent properties, text-based conditions often struggle to convey the detailed controls required for final generation results efficiently. This limitation becomes especially evident in certain scenarios where text prompts alone can not capture all necessary details, such as when depicting unique artistic styles or accurately rendering complex scenes. To this end, alongside text descriptions, a substantial body of research focuses on integrating novel conditional controls, such as user-drawn sketches and semantic masks, into T2I diffusion models (Qin et al., 2023; Ye et al., 2023; Zhang et al., 2023a; Mou et al., 2024). Despite investigations into the controllability of T2I diffusion models and their expanded applications (Gao et al., 2024; Li et al., 2024b; Xu et al., 2024; Gu et al., 2024), achieving precise and fine-grained control remains challenging. The primary issue lies in ensuring both the fidelity of the generated images and their semantic alignment with the provided conditions. To tackle this issue, some efforts, e.g., ControlNet++ (Li et al., 2024a) attempt to employ a pre-trained reward model to extract the corresponding condition from the generated images and enforce alignment between the specified condition and the output. However, we observe that the reward model inevitably produces inaccurate feedback (see Fig. 1). During the diffusion training, we add different levels of Guassian noise to the input, which increases the reconstruction challenge. It also leads to the distribution discrepancy between the generated image and the real image. As the timestep $t$ increases, such generation discrepancy increases. The reward model has not "seen" such generation discrepancy before, resulting false segmentation prediction, even if the generation is correctly aligned with the condition. Due to the diffusion compression process, we observe that even if $t=0$, there is some mis-alignment feedback as well. If we do not rectify such misleading rewards, enforcing alignment between the provided conditions and the inaccurate predictions on newly generated data will consistently compromise the training process.
@@ 2 Related Works
Conditional Generation. Recently, diffusion probabilistic models (Ho et al., 2020; Song et al., 2020a) have become an important cornerstone in general image generation. As the quality of generated images improves, a critical challenge remains, i.e., achieving precise control over these generative models to meet the intricate and varied demands of real-world applications adequately. Motivated by the development of guidance mechanism (Song et al., 2020b; Ho & Salimans, 2022), T2I diffusion models such as GLIDE (Nichol et al., 2021), Imagen (Saharia et al., 2022), DALL·E 2 (Ramesh et al., 2022) excel at modeling fine-grained structures and texture details. To facilitate the training of diffusion models on limited resources, Latent Diffusion Model (LDM) (Rombach et al., 2022) maps the diffusion process from pixel space to the latent space. Building upon the foundation of LDM framework, Stable Diffusion (SD) exhibits exceptional capabilities in T2I generation, gaining widespread usage within the community due to its open-source models. Despite the astonishing capabilities of T2I diffusion models, language, with its sparse and high-level semantic nature, is unsuitable for processing intricate and low-level control images, e.g., depth maps. To achieve conditional control in T2I diffusion models, researchers are exploring the integration of various control signals with text descriptions. ControlNet (Zhang et al., 2023a) integrates image-based conditions by incorporating an additional encoder copy into frozen T2I diffusion models via zero convolutions. This additional module links to the original UNet layers facilitates the integration of conditional inputs and prevent degradation in performance. Similarly, GLIGEN (Li et al., 2023) and T2I-Adapter (Mou et al., 2024) propose utilizing independent adapters (or extra modules) to synchronize internal knowledge within T2I models with external conditions. Furthermore, Cocktail (Hu et al., 2023) integrates multi-modal control information to yield more high-quality outputs. With the rapid development of large language models, some recent works have explored prompt engineering for regulated generation. For instance, ReCo (Zhang et al., 2023b) specifies text descriptions of different regions to model the fine-grained structure and aesthetic features. Control-GPT (Zhang et al., 2023b) applies GPT-4 (Achiam et al., 2023) as the sketch generator to incorporate control signals. However, one drawback is that their generated images usually deviate from input conditions. In this work, we take a closer look at denoised results and explicitly introduce the uncertainty to impose consistency constraints adaptively.
@@
Uncertainty-aware learning. As data-driven technologies rapidly advance, the imperative for model reliability grows. Measuring the "confidence" of predictions has been a long-standing problem. As a result, researchers are progressively focusing on uncertainty as a viable solution.  Kendall & Gal (2017) categorize uncertainty into two primary types: epistemic uncertainty and aleatoric uncertainty. The former epistemic uncertainty denotes model uncertainty, reflecting variations in model weights even when trained on the same dataset. The representative approach in this direction is Bayesian networks (Jensen & Nielsen, 2007; Neal, 2012), which focuses on learning the distribution of weights and estimates uncertainty by assessing the distribution variance. Similarly, Monte Carlo Dropout (Gal & Ghahramani, 2016) is proposed to randomly use varying dropout masks to simulate variational Bayesian approximation effectively. Another line of research on epistemic uncertainty (Raghu et al., 2019; Nandy et al., 2020; Lee & AlRegib, 2020) adapts the original model to measure prediction uncertainty directly. However, incorporating this additional uncertainty estimation typically compromises both the training efficiency and predictive accuracy of the model. In addition, ensemble methods (Lakshminarayanan et al., 2017; Malinin et al., 2019; Wenzel et al., 2020) combine various deterministic models in the prediction process to improve prediction accuracy, but is constrained by the computational burden associated with operating multiple independent networks and the requisite diversity across ensemble models. Another family of research tackles the inherent uncertainty in data, arising from noise interference or ambiguous annotations. This methodology for addressing uncertainty has been applied across various fields: such as person re-identification (Zhang et al., 2022), image retrieval (Chen et al., 2022), image classification (Litrico et al., 2023), hallucination detection (Zhang et al., 2024b) and 3D object detection (Zhang et al., 2024a). In particular,  Dou et al. (2022) effectively simulates uncertainty by utilizing model prediction variance to inject noise into the latent space. In a manner akin to (Kendall & Gal, 2017), a dynamic uncertainty-based loss is introduced to enhance training stability. Unlike existing uncertainty-based works, we explicitly introduce noise at varying intensities into the diffusion model to enable the quantification of uncertainty. It is worth noting that we explore the utilization of uncertainty in the untapped conditional image generation task, preserving the original diffusion learning and mitigating the adverse effects of inaccurate feedback from the reward model.
Figure 2: A brief overview of our pipeline. Here, we take the segmentation mask as a conditional generation example. (a) Conditional Generation. Given text, source image $x_{0}$, and the conditional control $c$, we extract feature $z_{0}$, $f_{t}$, $f_{c}$, respectively. Then, we fine-tune the Diffusion model to generate two intermediate features for the image decoder. (b) Uncertainty Learning. Given the two features, we decode the two images, i.e., $\hat{x}_{0}^{1}$ and $\hat{x}_{0}^{2}$. Then we apply the reward model to obtain the two layout predictions $\hat{c}_{1}$ and $\hat{c}_{2}$. We leverage the KL-divergence prediction discrepancy between $\hat{c}_{1},\hat{c}_{2}$ as the uncertainty indicator $U_{1},U_{2}$ (see Eq. 2). Based on $U_{1},U_{2}$, we then calculate the rectified reward loss between the predicted label $\hat{c}_{1},\hat{c}_{2}$ and the ground-truth label $c$, as Eq. 3.
@@ 3.1 Uncertainty Estimation
$z_{1}=\sqrt{\bar{\alpha}_{t_{1}}}z_{0}+\sqrt{1-\bar{\alpha}_{t_{1}}}\epsilon,\quad z_{2}=\sqrt{\bar{\alpha}_{t_{2}}}z_{0}+\sqrt{1-\bar{\alpha}_{t_{2}}}\epsilon,\quad\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\vspace{-.1in}$ (1)
Following the ControlNet pipeline (Zhang et al., 2023b), we further fuse text condition $f_{t}$ and image condition $f_{c}$ to predict the added noise. After removing the predicted noise, we obtain the recovered latent, i.e., $\hat{z}_{0}^{1}$ and $\hat{z}_{0}^{2}$. Then, given the latent $\hat{z}_{0}^{1}$ and $\hat{z}_{0}^{2}$, the pretrained decoder is to reconstruct the input image as $\hat{x}_{0}^{1}$ and $\hat{x}_{0}^{2}$, respectively. To align the condition within the generated images, we follow the existing works and apply an off-the-shelf reward model $D$ to quantify the consistency between the input condition $c$ and the corresponding output condition of the generated images. However, as shown in Fig. 1(b), the reward model usually contains inaccurate feedback, even if the visual layout has already aligned with the condition. If we consistently apply strong supervision on the consistency between inaccurate rewards and conditions predicted on the newly generated data, the backpropagation of incorrect gradients will significantly compromise the model. To estimate inaccrate rewards, we explicitly leverage two diffusion forwards for the identical input conditions. We compare the reward discrepancy between extracted conditions $\hat{c}_{1},\hat{c}_{2}$ from generated images, which can be considered as a reward indicator at the current timestep. For the segmentation map condition, considering the probability output, we quantify uncertainty by calculating the KL-divergence between the extracted conditions of two generations at the pixel level as:
@@ 3.2 Uncertainty Regularization
The existing controllability modeling methods (Li et al., 2024a) usually adopt consistency loss between the input condition and the extracted condition, which can be regarded as a pixel-wise supervision. For example, when using segmentation mask as input condition, $\mathcal{L}^{c}$ can be defined as a per-pixel cross-entropy loss as:
@@
$\mu=\begin{cases}\mu_{0}&\text{if }t\leq t_{\text{thre}}\\ 0&\text{if }t>t_{\text{thre}}\end{cases}$ (7)
where $\mu_{0}$ is the consistency weight. If $t>t_{\text{thre}}$, we only consider the diffusion optimization as existing works (Zhao et al., 2023a). if $t\leq t_{\text{thre}}$, we incorporate uncertainty-aware reward modeling.
Discussion. 1). Why set timestep threshold $t_{\text{thre}}$ for $\mu$? Considering that when $t$, i.e., $t_{1}$ and $t_{2}$, is large, $z_{t}$ approaches the random noise $\epsilon$, leading to more diverse recovered $\hat{x}_{0}$ (see Fig. 1). In such cases, it becomes difficult to ensure that two outputs with identical conditions remain close, which is necessary for our uncertainty estimation, as it requires visually similar inputs to evaluate the reward model. In practise, we, therefore, empirically set a threshold to avoid the large timesteps for the reward modeling. We have conducted the experiment on the choice of $t_{\text{thre}}$ (see Section 4.3). 2). What is the advantage of uncertainty regularization in the reward feedback? The advantage of uncertainty regularization in the reward feedback is that it helps to mitigate the adverse effects of imprecise feedback from the reward model. Specifically, by incorporating uncertainty estimation, the proposed method, Ctrl-U, can adaptively adjust the weights of the rewards during training. Rewards with lower uncertainty are given higher loss weights, ensuring that more reliable feedback has a stronger influence on the training process. Conversely, rewards with higher uncertainty are given reduced weights, allowing for greater variability and preventing the model from being overly influenced by potentially inaccurate feedback. This adaptive rectification of rewards based on their uncertainty improves the consistency and reliability of the training process. As a result, the model becomes more robust to the diverse and sometimes unpredictable nature of newly generated data, leading to better controllability and generation quality.